A sample of air at 2.50 atm is cooled from 450 K to 290 K. If the volume remains constant, what is the final pressure?
1
Expert's answer
2018-05-02T07:56:12-0400
P1/T1=P2/T2
P2=P1*T2/T1
P2=2.5atm*290K/450K=1.61atm